# VisJudge-Bench: Aesthetics and Quality Assessment of Visualizations

### v2

@@ C.1.3 Dashboard Evaluation Criteria
Prompt: Dashboard - Data Fidelity Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the dashboard faithfully represents business data in each component and in the overall composition. Focus on appropriate axes/baselines, reasonable scale ranges, consistency between encodings and numeric labels, avoidance of aspect-ratio deformation, cropping/stretching, 3D distortion, or improper broken axes; also consider consistency of comparable metrics across components. Scoring criteria: S P 1. Severe misrepresentation or impossible to judge due to missing/unreadable key elements in important components, or clear distortions (e.g., improper bar baselines, distorted 3D, obvious deformation), undermining data fidelity. S 2. Important issues likely to mislead in one or more components: inappropriate baselines/scales, inconsistent encoding vs labels, selective ranges, or missing key elements causing notable uncertainty at the dashboard level. S 3. Basically acceptable overall: minor issues or uncertainties in parts, but no obvious distortion; most components use reasonable scales/encodings consistent with labels. S 4. Accurate and standardized: components and overall dashboard use appropriate axes/baselines and reasonable scales; encodings match labels; no visible distortion. S 5. Highly faithful: all components present complete and clear elements; axes/baselines and scales are well-justified; encodings and labels fully consistent; no cropping/deformation/3D misuse; consistent comparability across related components.
@@
Prompt: Dashboard - Semantic Readability Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Based on complete and clear dashboard elements, evaluates whether users can understand the meaning of these elements and the information the dashboard conveys, including clarity of visual encoding (whether meanings of colors, shapes, sizes, gauge pointers, status indicators are clear) and clarity of information communication (whether users can accurately understand various parts and overall business meaning). Scoring criteria: S 1. Dashboard element meanings are completely unclear, cannot understand what charts, indicators, color encoding represent in business terms, users cannot obtain meaningful information. S 2. Dashboard element meanings are vague and unclear, insufficient business indicator explanations, unclear color encoding and status indication, users have difficulty understanding and need extensive guessing to understand content. S 3. Dashboard element meanings are basically clear, users can understand main business information and indicator meanings, but still have understanding barriers in some encoding or labeling aspects. S 4. Dashboard element meanings are clear and definite, business indicators have appropriate explanations, clear visual encoding, users can smoothly understand the business information conveyed by the dashboard. S 5. Dashboard element meanings are completely clear, business indicator explanations are detailed and accurate, all visual encoding has clear interpretation, users can understand all business information in the dashboard without any barriers.
@@
Prompt: Dashboard - Insight Discovery Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the entire dashboard can easily identify significant and business-valuable key insights, whether key business indicators (KPIs) prominently display important information, and whether it can quickly discover business issues or opportunities that require attention. Scoring criteria: S 1. Completely unable to provide any valuable business insights, key indicators not prominent, no significant or meaningful business discoveries, insight delivery completely failed. S 2. Difficult to identify meaningful business insights, insufficient highlighting of key indicators, most presented business information is trivial or obvious, lacking decision value. S 3. Can identify some basic business insights, key indicators have some prominence, but the significance and business value of insights are average, with limited guiding effects. S 4. Can clearly identify obvious and useful business insights, key indicators prominently displayed, discovered business patterns or anomalies have certain practical value and decision reference significance. S 5. Very intuitively reveals profound and highly business-valuable insights, key indicators extremely prominent, can quickly discover significant business trends, important anomalies, or key opportunities with important guiding significance for business decisions.
@@
Prompt: Dashboard - Design Style Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the overall design of the dashboard has innovation and uniqueness, whether it can show attractive visual effects and business professionalism through novel design elements, unique style characteristics, and professional design language. Scoring criteria: S 1. Dashboard design completely lacks innovation, style is outdated or chaotic, no uniqueness, seriously insufficient professionalism. S 2. Dashboard design lacks innovation, style is quite ordinary, basically uses common design techniques, insufficient uniqueness and professionalism. S 3. Dashboard design has some innovative elements, but overall is relatively conventional, style characteristics not prominent enough, average professionalism. S 4. Dashboard design has strong innovation, style is quite unique and professional, with certain originality and good business sense. S 5. Dashboard design is extremely innovative, style is unique, novel, and professional, overall presents strong originality and high business aesthetics, leaving deep impressions.
@@
Prompt: Dashboard - Visual Composition Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether the spatial layout of dashboard components and overall arrangement are reasonable and beautiful, including size proportions of various blocks, alignment relationships between components, spacing distribution, information density control, etc., forming a harmonious, orderly, and beautiful visual whole. Scoring criteria: S 1. Component layouts are chaotic or overall arrangement is seriously unreasonable, size proportions severely unbalanced, spacing distribution is disorderly, poor information density handling. S 2. Component layouts or overall arrangement have obvious unreasonable aspects, inappropriate size proportions, uneven spacing distribution, or information too crowded. S 3. Component layouts are basically reasonable, overall arrangement is acceptable, but may have small issues in proportional relationships, spacing handling, or information density. S 4. Component layouts are good, overall arrangement is reasonable and beautiful, appropriate size proportions, proper spacing distribution, good information density control. S 5. Component layouts are perfect, overall arrangement is extremely reasonable and beautiful, coordinated size proportions, aesthetic spacing distribution, appropriate information density control, extremely high space utilization efficiency.
@@
Prompt: Dashboard - Color Harmony Evaluation Questions and Scoring Criteria (1-5 Scale) Description: Evaluates whether color choices in the dashboard are appropriate and beautiful, including color combinations within various components, coordination of overall color schemes, unity of tone, and appropriate control of color quantity and saturation, ensuring both beauty and compliance with business professional requirements. Scoring criteria: S 1. Component color choices are very inappropriate, overall color matching is chaotic, tone conflicts or color use is seriously inappropriate, very poor aesthetic effect. S 2. Component color choices are not appropriate enough, overall color matching is not coordinated enough, tone not unified or color use inappropriate, affecting aesthetic effect. S 3. Component color choices are basically appropriate, overall color matching is acceptable, but may have small issues in tone unity, saturation, or business sense. S 4. Component color choices are relatively appropriate, overall color schemes are quite coordinated, tone basically unified, appropriate color use, good business sense. S 5. Component color choices are very appropriate, overall color schemes are highly coordinated and unified, perfect tone matching, appropriate control of color quantity and saturation, presenting excellent business aesthetics.
@@ D.2 Reward Function
This component measures the proximity between the model’s predicted scores across the six dimensions and the average score, and the human-annotated ground-truth values. We employ a smooth exponential decay function to calculate the reward for each individual score:
@@ Appendix E Extended Experiments
E.1 Generalization to Real-World Applications
@@
E.1.1 Experimental Setup
MatPlotAgent (Visualization Generation). We integrate VisJudge into MatPlotAgent (Yang et al., 2024) as a feedback mechanism for iterative quality improvement. We test seven generation models (GPT-5, GPT-4o, Claude-4-Sonnet, Claude-3.5-Sonnet, Gemini-2.5-Pro, Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-7B-Instruct), comparing baseline (no feedback) against feedback from 7B base model, 72B base model, and VisJudge. Final visualizations are automatically evaluated using GPT-4o on a 0-100 scale, following the original MatPlotAgent evaluation setup, where higher scores indicate better overall visualization quality.
@@ E.1.2 Results and Analysis
E.1.3 Cross-Architecture Generalization

### v3

@@ C.1.3 Dashboard Evaluation Criteria
Prompt: Dashboard - Data Fidelity Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the dashboard faithfully represents business data in each component and in the overall composition. Focus on appropriate axes/baselines, reasonable scale ranges, consistency between encodings and numeric labels, avoidance of aspect-ratio deformation, cropping/stretching, 3D distortion, or improper broken axes; also consider consistency of comparable metrics across components. Scoring criteria: S 1. Severe misrepresentation or impossible to judge due to missing/unreadable key elements in important components, or clear distortions (e.g., improper bar baselines, distorted 3D, obvious deformation), undermining data fidelity. S 2. Important issues likely to mislead in one or more components: inappropriate baselines/scales, inconsistent encoding vs labels, selective ranges, or missing key elements causing notable uncertainty at the dashboard level. S 3. Basically acceptable overall: minor issues or uncertainties in parts, but no obvious distortion; most components use reasonable scales/encodings consistent with labels. S 4. Accurate and standardized: components and overall dashboard use appropriate axes/baselines and reasonable scales; encodings match labels; no visible distortion. S 5. Highly faithful: all components present complete and clear elements; axes/baselines and scales are well-justified; encodings and labels fully consistent; no cropping/deformation/3D misuse; consistent comparability across related components.
@@
Prompt: Dashboard - Semantic Readability Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Based on complete and clear dashboard elements, evaluates whether users can understand the meaning of these elements and the information the dashboard conveys, including clarity of visual encoding (whether meanings of colors, shapes, sizes, gauge pointers, status indicators are clear) and clarity of information communication (whether users can accurately understand various parts and overall business meaning). Scoring criteria: S 1. Dashboard element meanings are completely unclear, cannot understand what charts, indicators, color encoding represent in business terms, users cannot obtain meaningful information. S 2. Dashboard element meanings are vague and unclear, insufficient business indicator explanations, unclear color encoding and status indication, users have difficulty understanding and need extensive guessing to understand content. S 3. Dashboard element meanings are basically clear, users can understand main business information and indicator meanings, but still have understanding barriers in some encoding or labeling aspects. S 4. Dashboard element meanings are clear and definite, business indicators have appropriate explanations, clear visual encoding, users can smoothly understand the business information conveyed by the dashboard. S 5. Dashboard element meanings are completely clear, business indicator explanations are detailed and accurate, all visual encoding has clear interpretation, users can understand all business information in the dashboard without any barriers.
@@
Prompt: Dashboard - Insight Discovery Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the entire dashboard can easily identify significant and business-valuable key insights, whether key business indicators (KPIs) prominently display important information, and whether it can quickly discover business issues or opportunities that require attention. Scoring criteria: S 1. Completely unable to provide any valuable business insights, key indicators not prominent, no significant or meaningful business discoveries, insight delivery completely failed. S 2. Difficult to identify meaningful business insights, insufficient highlighting of key indicators, most presented business information is trivial or obvious, lacking decision value. S 3. Can identify some basic business insights, key indicators have some prominence, but the significance and business value of insights are average, with limited guiding effects. S 4. Can clearly identify obvious and useful business insights, key indicators prominently displayed, discovered business patterns or anomalies have certain practical value and decision reference significance. S 5. Very intuitively reveals profound and highly business-valuable insights, key indicators extremely prominent, can quickly discover significant business trends, important anomalies, or key opportunities with important guiding significance for business decisions.
@@
Prompt: Dashboard - Design Style Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the overall design of the dashboard has innovation and uniqueness, whether it can show attractive visual effects and business professionalism through novel design elements, unique style characteristics, and professional design language. Scoring criteria: S 1. Dashboard design completely lacks innovation, style is outdated or chaotic, no uniqueness, seriously insufficient professionalism. S 2. Dashboard design lacks innovation, style is quite ordinary, basically uses common design techniques, insufficient uniqueness and professionalism. S 3. Dashboard design has some innovative elements, but overall is relatively conventional, style characteristics not prominent enough, average professionalism. S 4. Dashboard design has strong innovation, style is quite unique and professional, with certain originality and good business sense. S 5. Dashboard design is extremely innovative, style is unique, novel, and professional, overall presents strong originality and high business aesthetics, leaving deep impressions.
@@
Prompt: Dashboard - Visual Composition Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether the spatial layout of dashboard components and overall arrangement are reasonable and beautiful, including size proportions of various blocks, alignment relationships between components, spacing distribution, information density control, etc., forming a harmonious, orderly, and beautiful visual whole. Scoring criteria: S 1. Component layouts are chaotic or overall arrangement is seriously unreasonable, size proportions severely unbalanced, spacing distribution is disorderly, poor information density handling. S 2. Component layouts or overall arrangement have obvious unreasonable aspects, inappropriate size proportions, uneven spacing distribution, or information too crowded. S 3. Component layouts are basically reasonable, overall arrangement is acceptable, but may have small issues in proportional relationships, spacing handling, or information density. S 4. Component layouts are good, overall arrangement is reasonable and beautiful, appropriate size proportions, proper spacing distribution, good information density control. S 5. Component layouts are perfect, overall arrangement is extremely reasonable and beautiful, coordinated size proportions, aesthetic spacing distribution, appropriate information density control, extremely high space utilization efficiency.
@@
Prompt: Dashboard - Color Harmony Evaluation Questions and Scoring Criteria (1–5 Scale) Description: Evaluates whether color choices in the dashboard are appropriate and beautiful, including color combinations within various components, coordination of overall color schemes, unity of tone, and appropriate control of color quantity and saturation, ensuring both beauty and compliance with business professional requirements. Scoring criteria: S 1. Component color choices are very inappropriate, overall color matching is chaotic, tone conflicts or color use is seriously inappropriate, very poor aesthetic effect. S 2. Component color choices are not appropriate enough, overall color matching is not coordinated enough, tone not unified or color use inappropriate, affecting aesthetic effect. S 3. Component color choices are basically appropriate, overall color matching is acceptable, but may have small issues in tone unity, saturation, or business sense. S 4. Component color choices are relatively appropriate, overall color schemes are quite coordinated, tone basically unified, appropriate color use, good business sense. S 5. Component color choices are very appropriate, overall color schemes are highly coordinated and unified, perfect tone matching, appropriate control of color quantity and saturation, presenting excellent business aesthetics.
@@ D.2 Reward Function
This component measures the proximity between the model’s predicted scores (across the six dimensions and the overall average score) and the human-annotated ground-truth values. We employ a smooth exponential decay function to calculate the reward for each individual score:
@@ Appendix E Extended Experiments
(a) Training data scale vs. human-model correlation
(b) Training data scale vs. prediction error
E.1 How Does Training Data Scale Affect Model Performance?
E.2 Generalization to Real-World Applications
@@
E.2.1 Experimental Setup
MatPlotAgent (Visualization Generation). We integrate VisJudge into MatPlotAgent (Yang et al., 2024) as a feedback mechanism for iterative quality improvement. We test seven generation models (GPT-5, GPT-4o, Claude-4-Sonnet, Claude-3.5-Sonnet, Gemini-2.5-Pro, Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-7B-Instruct), comparing baseline (no feedback) against feedback from 7B base model, 72B base model, and VisJudge. Final visualizations are automatically evaluated using GPT-4o on a 0–100 scale, following the original MatPlotAgent evaluation setup, where higher scores indicate better overall visualization quality.
@@ E.2.2 Results and Analysis
E.2.3 Cross-Architecture Generalization